Let demand be given by Q = 150 - P + 2Y. This is the same for all problems of this type. Let r = 10%. Let Y = 50 across both periods. Let MC = 0. Let reserves = 200. Consider the basic two-period model. What is consumption of the resource in the present IF THE OWNER OF THE RESOURCE IS A MONOPOLIST?

94.29

98.81

101.19

105.71

none of the above
Answer #1

Answer option 3)

Q= 150-P+2*50

Q= 250-P, MC= 0, r= .1,

Q1+Q2= 200

As P= 250-Q

MR= 250-2Q

.

so, at MONOPOLY extraction

MR(q1) - MC = [ MR(q2) - MC]/(1+r)

250-2Q1 = (1/1+r )*(250-2Q2)

250-2Q1 = (250-2Q2)/1.1

275- 2.2*Q1 = (250-2*Q2)

Put Q2= 200-Q1

25-2.2*Q1 = 2(Q1-200)

425= 4.2Q1

Q1*=101.19